# JWST imaging of the Pleiades: anisotropy of turbulence in the cold neutral medium

### v1

@@ Appendix A Obtaining images of the diffuse emission
This appendix details and illustrates the methods used to create maps of diffuse emission out of the JWST mosaics. Fig. 11 presents a zoomed-in view of the data to illustrate the diversity of emission sources we need to subtract or mask to obtain maps that can be used to statistically analyze the diffuse emission.
@@ A.1 Removing diffraction spikes from off-field point sources
The bright star Merope in the vicinity of our field of view produces three diffraction spikes that are bright features across Field 1 (see bottom left panel of Fig. 1). The central one to the East of Merope aligns with the NIRCam detector arrays, and the other two are at . The Merope spike with the lowest declination extends to Field 2. We model and remove the three spikes from Merope in Field 1, the spike from Merope that extends to Field 2, and two fainter spikes from PQ Tau in Field 1. The first three spikes are clearly visible on the bottom right panel in Fig. 1, while the fourth is barely noticeable on the bottom left panel. The last two are too faint to be visible in the Figure. We illustrate our modeling with the eastern spike.
To model the oblique spikes, we rotate the JWST mosaic in array coordinates by plus or minus 60 degrees to apply the same procedure.
@@
We cut a rectangular section of the observations surrounding the diffraction spike, which is 75 to 175 pixels wide on each side (see the first panel in Fig. 12). The width chosen depends on the intensity of the spike. In the following, columns refer to one-dimensional cuts within the cutout image, across the spike, and rows to one-dimensional cuts that are parallel to the spike. The steps involved in modeling and subtracting the spikes are illustrated in Fig. 12.
@@
The first step is to estimate the background emission around the spike. To achieve this, we mask the inner 50 to 150 pixels where the spike dominates the emission. We also mask any value above from the median in each column (using sigma_clip from Astropy). The data along each column are fitted with a third-order polynomial function, which we use to interpolate and fill in the diffuse emission at the position of the masked pixels. We apply a median filter on the polynomial interpolation, with a width along the spike orientation of 100 pixels, to filter out irregularities due to compact sources. The resulting map is shown in the second panel of Fig. 12. We subtract this background from the initial cutout image, which yields a map of the spike. This map still contains point and compact sources, as well as some residual diffuse emission (see the third panel in Fig. 12).
@@
The second step involves modeling the spike. We estimate the gradient along the diffraction spike, assuming that it can be modeled by one polynomial function that applies to each row of pixels. To obtain this polynomial function, we calculate the -clipped median of each column across the width of the spike: 150 pixels for the eastern spike of Merope, 100 pixels for the other spikes. This gives us the typical longitudinal gradient profile of the spike, which we fit with a 7th order polynomial function (see Fig. 13).
@@
Finally, we normalize the first approximation of the spike we obtained earlier by dividing it, for each column across the spike, by the polynomial values of the longitudinal gradient. For each column, we then select the nearest 100 columns to compute the -clipped median of this selection for each row. This gives us the profiles across the spike, assuming it does not vary on scales much shorter than 100 pixels or 6.3″(see Fig. 14).
@@
We note that the pattern varies smoothly along the spike. In particular, we see the peak of the profile shifts by about 2 pixels over 2000 pixels. A larger shift seems to occur at the faintest end of the spike, far from Merope, where the relative intensity of the spike is the lowest. However, even there, the shift is not greater than 7 pixels compared to the other end of the spike. Next, we rescale the normalized profiles with the polynomial gradient to obtain the spike model shown in the fourth panel of Fig. 12. Finally, we subtract the model from the initial map (see the fifth panel in Fig. 12). To assess the accuracy with which the spike is removed, we compared the residual background in the spike image, away from the central area, relative to the initial map. We find that, in the case of Merope’s straight spike, the residual is typically lower than 1% at more than 75 pixels (4.7 ″) from the center of the spike (see Fig. 15).
@@
A limitation in our method is visible in Fig. 16. Closer to Merope, the pattern across the spike varies on scales smaller than 100 pixels (or 6.1″), leaving a residual trace near the edge of our observations. However, the alternative (i.e. using a smaller window than 100 pixels) would not properly filter out the sources “hidden” by the spikes. We compare these residuals with local dust emission by normalizing the profiles at the site of the spike to their median values (see Figure 17). In most cases, the residual of the spike is barely noticeable above the intrinsic variations of the diffuse emission, at a level of 5% of the median value of the local background. However, near Merope, the residual reaches more than 10% of that level but remains mostly random.
@@
At this point, we thus have four mosaics, two for each field, with all large diffraction spikes removed. For each pair of mosaics, we remove the glitches by comparing both mosaics. Each pixel whose value varies by more than 0.5 MJy/sr is replaced by the lower value in the pair.
A.2 Multiscale background
@@
Once the large spikes have been removed from the data, we model the diffuse emission. We used mosaics of Fields 1 and 2 from which the spikes have been removed. We compute their global median by clipping pixels above 2- and subtract the result from the images.
@@
Next, we subtract the different scales of the diffuse emission by repeating three times the following steps: (1) we compute the standard deviation of the 2--clipped background subtracted image, (2) we mask the bright pixels whose values are more than 15- above the median of the entire image, (3) we perform a 2D median-filtering with a 4- clipping without the masked pixels, and (4) we subtract the median-filtered image. We use the Background2D class from photutils to perform these steps (Bradley et al. 2023). We apply the 2D median filter using rectangular boxes elongated along the direction of the detector arrays, which is close to the orientation of the striations. We use boxes with mean sizes of 100, 40 and 25 pixels. The ratio between the long and short sizes is 2 at all three scales. Figure 18 illustrates the composition of the background map as smaller scales are included, as well as the residual map obtained after background subtraction for a sub-image of Field 1. There is structure in the diffuse emission at scales smaller than 25 pixels (i.e. 1.5 arcsec). This contributes noticeably, albeit minimally, to the residuals, even in Field 1, where the diffuse emission is strongest.
@@ A.3 Removing point sources within the field
After subtracting the multiscale background, we use the residual images to identify point sources and model the PSF. We identify point sources by recursively selecting the brightest pixel in the four combined mosaics. For each such pixel, we look for all other neighbor bright pixels and label the cluster as a source before looking for the next brightest pixel in the observations. We extract in this way the brightest 400 sources from our observations. We then visually inspect them all to separate point sources from extended sources, and end up with 175 point sources. We do not use publicly available source extraction algorithms as they tend to identify multiple sources within each JWST PSF, as those are highly structured for the brightest sources. Our method is more robust and still functional given the small number of sources we want to extract.
To build an empirical PSF (ePSF), we follow the method developed by Anderson and King (2000) and Anderson (2016), which is implemented as the EPSFBuilder class in photutils, adapting it to our specific needs. In particular, we want to clean the spikes of the point sources better than the cores because the former contaminate the diffuse emission more than the latter.
To combine the images of the selected point sources, we first center each image to subpixel accuracy using the theoretical PSF (tPSF) of WebbPSF (Perrin et al. 2014) on a 301 by 301 pixel ( by ). Each PSF is fitted in a squared annulus region, instead of its core, and we find that using a region of 45 by 45 pixels, with the 9 by 9 innermost pixels masked, gives the best results. The difference between the position of the brightest pixel and the computed centers of the PSFs is typically less than 0.4 pixel.
The centered PSFs are then stacked in an iterative way by our modified EPSFBuilder to generate the ePSF for our observations. Here, we use the same square annulus as before to align the individual PSFs with each other. The method also returns the individual best fit for each source in our collection, in their 301 by 301 original pixel grid, which we subtract from the observations, except for the six brightest point sources that have diffraction spikes extending beyond 20″ (see Fig. 19). There are four point sources in Field 1 and two in Field 2, for which the core is highly saturated. If we had removed their ePSF best fit, this would have led to significant residuals across the entire 301 by 301 pixel region. For these sources, we instead wait until we add back the multi-scale background, after which we remove their diffraction spikes in the same way we use for the stars off field (see Sect. A.1).
@@ A.4 Masking the remaining sources
At this stage, we have removed all diffraction spikes from sources within and without our field of observations. We are left with masking the saturated core of bright point sources and compact sources.
To build the mask of the remaining sources, we first identify pixels that correspond to significant residuals after subtracting the ePSF fit. We compute the 2--clipped standard deviation for each 301 by 301 pixel image of a point source, keeping only values above a floor of 0.1 MJy/sr. Any pixel in the image whose value in the residual of the ePSF fit is greater than 10 times this standard deviation is flagged. Any pixel whose value in the ePSF fit is greater than 15 times this standard deviation is also flagged. These pixels represent less than 0.5% of all pixels in the four sub-mosaics and are the first included in a mask.
Next, we mask additional bright pixels. We start by creating a multiscale background in an identical manner to what was done in the previous subsection, but now starting from a cleaner map. We then select all pixels where the residual emission after subtraction of the multiscale background is larger than a brightness threshold (0.035 MJy/sr for the fiducial mask). We then identify small and large clusters of masked pixels: those with fewer than or more than nine connected masked pixels. We verify that the small clusters typically correspond to faint stars or isolated glitches. Large clusters correspond either to the core of galaxies, whose shapes and sizes are very diverse (see Figure 11), or to the few bright point sources for which the ePSF method was not applied. Since the brightness threshold leaves the edges of bright sources unmasked, we extend the large clusters by a buffer zone of fifteen additional pixels. To do this, we use the skimage.morphology package (van der Walt et al. 2014). Finally, we add back the small clusters of pixels to the mask.
Starting from mosaics containing large diffraction spikes, point sources, and extended sources, we have step by step removed the spikes from both the off-field and in-field sources and masked both the core of the point sources and the extended sources. This leaves us with mosaics containing only diffuse emission, which are used to perform the Fourier analysis presented in Sect. 3.
@@ Appendix B Magnetized interstellar matter near the Pleiades stars
Notes. (a) Dust temperature from Planck data at 5’ resolution. (b) Dust opacity at GHz from Planck data at 5’ resolution. (c) Stokes , and at 353 GHz measured at the center of the JWST images on Planck maps smoothed to a beam. The values for Fields 1 and 2 are background subtracted. (d) Brightness at m measured at the center of the JWST images on the WISE all-sky map (Meisner and Finkbeiner 2014). The values for Fields 1 and 2 are background subtracted. (e) Median values over the background area. The error bars are the standard deviation of pixel values. Uncertainties on the median values are one order of magnitude smaller.
@@
We used the dust temperature and opacity maps derived from fitting the dust spectral energy distribution at far-infrared wavelengths to a modified black body (Planck Collaboration 2016b), and the dust polarization maps at 353 GHz (Planck Collaboration 2020b). We also used the full sky image produced by Meisner and Finkbeiner (2014) from the WISE data at m888Image downloaded from https://faun.rc.fas.harvard.edu/ameisner/wssa/healpix.html. The angular resolution of these maps is 5’. However, Stokes and are smoothed to 7’ in order to improve the signal-to-noise ratio of the dust polarization.
The Planck and WISE maps offer complementary insights into the interstellar medium. The dust temperature traces the localized heating of dust grains by the Pleiades stars, while the dust opacity indicates the gas column density. The conversion factor from to for translucent molecular gas is (Planck Collaboration 2014). The dust polarization angles are perpendicular to the magnetic field component in the plane of the sky. The WISE image traces the emission of PAHs, the brightness of which scales linearly with the product of the intensity of the radiation field and the gas column density (Draine and Li 2007; Compiègne et al. 2011). Measurements of these observational tracers computed at the center positions of the four JWST images and for the surrounding background are listed in Table 7. The background measurements correspond to the median value calculated for pixels within of Merope where is smaller than to exclude the Pleiades area.
Figure 20 shows the Planck maps of dust temperature and submillimeter opacity map at GHz overlaid on the m WISE image at 12″resolution999The image was downloaded from https://irsa.ipac.caltech.edu/applications/wise. At the distance of the Pleiades stars, the angular size of the image corresponds to 3 pc. The dust temperature exceeds the background value across the full extent of the mid-infrared emission from the nebula. The opacity contours delineate an interstellar cloud to the south of Merope. It corresponds to a small cloud identified by Ungerechts and Thaddeus (1987) in their survey with a peak CO(1-0) brightness of . The Merope star and the two JWST fields are located at the edge of this translucent cloud. The overlap between the contours of and indicates that the cloud is locally heated by Merope.
@@ Appendix C Association with the cold neutral medium
In the paper introduction, we claim that the JWST Pleiades observations probe the structure of the CNM within a few parsecs from Merope. To support and discuss this statement, we use the Planck and WISE data presented in Appendix B. These data provide a global perspective on the JWST observations. The angular resolution of the Planck data corresponds to a scale of 0.2 pc at the distance of Merope.
@@
Merope lies to the front of the large (150 pc in diameter) Per-Tau bubble that was identified by Bialy et al. (2021) using 3 dimensional stellar reddening data. The star is close to the outer surface of the shell that surrounds the bubble. On smaller scales, the dust emission data presented in Appendix B show that Merope and the two JWST fields are located at the edge of a translucent cloud, with a few parsecs diameter. Table 7 lists the values of the dust temperature and the dust opacity in the center positions of the four JWST images, as well as for the surrounding background. The values of the Planck Stokes parameters at 353 GHz, and the PAH emission at m, measured on the WISE map of the entire sky produced by Meisner and Finkbeiner (2014), are also listed in the table. We used these data to estimate the radiation field intensity and gas column density at the location of the JWST fields.
@@
The radiation field intensity may be estimated from the dust temperature using the formula:
@@
with a spectral index (Planck Collaboration 2014). The factor measures the radiation field intensity normalized to its value within the background ISM, away from the Pleiades stars. This formula yields values ranging from 6 to 9 at the position of the JWST fields. The opacity values listed in Table 7 include the Galactic background. To estimate the column density of the cloud associated with Merope, we need to subtract this contribution. This subtraction introduces a large uncertainty equal to the dispersion of opacity values measured over the background area, which is listed as the error bar in the background line of the table. Taking into account this uncertainty, we consider that the column density of the cloud at the positions of the JWST fields cannot be much larger than .
@@
The m brightness allows os to crosscheck the estimates of and . In the solar neighborhood, the mean emission at m is for a column density of (Compiègne et al. 2011). For fixed PAH abundance and emission properties, this emission scales linearly with the product (Draine and Li 2007; Compiègne et al. 2011). Based on the m brightness values in Table 7, we estimate the product to be 16 and 7 for Fields 1 and 2, where is expressed in units of . These values suggest that the estimates of G and derived from dust temperature and opacity are somewhat underestimated. This could be due to beam smoothing and the superposition of cloud and background emissions, which were not considered when fitting the dust temperature and opacity. In any case, the estimates of and support our assertion that the JWST observations characterize emission from CNM within a few parsecs from Merope. The value of is 10 at a distance of 0.5 pc from Merope. The contribution to of warm gas with densities of a few tenths across this distance is negligible.
@@
We also estimate the contribution of the Galactic background to the JWST power spectra. To achieve this, we use the scaling law established by Miville-Deschênes et al. (2007) when analyzing IRAS data. The fraction of power contributed by the Galactic background is estimated using the following formula:
@@
where , the brightness in the center of a JWST image in the F335M filter, is obtained by multiplying the values and in Table 1. To obtain the background brightness in the F335M filter, we scaled the background brightness at m listed in Table 7 by the intensity ratio , as determined in the study of Flagey et al. (2006). We obtain values of % for the Field 1 images and between one and a few % for the Field 2 images. These low values support the link that the paper makes between the JWST power spectra and the CNM near the Pleiades stars.

### v2

@@ Appendix A Obtaining images of the diffuse emission
This appendix details and illustrates the methods used to create maps of diffuse emission out of the JWST mosaics. Figure 11 presents a zoomed-in view of the data to illustrate the diversity of emission sources we needed to subtract or mask to obtain maps that can be used to statistically analyze the diffuse emission.
@@ A.1 Removing diffraction spikes from off-field point sources
The bright star Merope in the vicinity of our field of view produces three diffraction spikes that are bright features across Field 1 (see the bottom-left panel of Fig. 1). The central one to the east of Merope aligns with the NIRCam detector arrays, and the other two are at . The Merope spike with the lowest declination extends to Field 2. We modeled and removed the three spikes from Merope in Field 1, the spike from Merope that extends to Field 2, and two fainter spikes from PQ Tau in Field 1. The first three spikes are clearly visible on the bottom right panel in Fig. 1, while the fourth is barely noticeable on the bottom left panel. The last two are too faint to be visible in the figure. We illustrate our modeling with the eastern spike. To model the oblique spikes, we rotated the JWST mosaic in array coordinates by plus or minus 60 degrees to apply the same procedure.
@@
We cut a rectangular section of the observations surrounding the diffraction spike, which is 75 to 175 pixels wide on each side (see the first panel in Fig. 12). The width chosen depends on the intensity of the spike. In the following, columns refer to 1D cuts within the cutout image, across the spike, and rows to 1D cuts that are parallel to the spike. The steps involved in modeling and subtracting the spikes are illustrated in Fig. 12.
@@
The first step is to estimate the background emission around the spike. To achieve this, we masked the inner 50 to 150 pixels where the spike dominates the emission. We also masked any value above from the median in each column (using sigma_clip from Astropy). The data along each column are fitted with a third-order polynomial function, which we used to interpolate and fill in the diffuse emission at the position of the masked pixels. We applied a median filter to the polynomial interpolation, with a width along the spike orientation of 100 pixels, to filter out irregularities due to compact sources. The resulting map is shown in the second panel of Fig. 12. We subtracted this background from the initial cutout image, which yielded a map of the spike. This map still contains point and compact sources, as well as some residual diffuse emission (see the third panel in Fig. 12).
@@
The second step involves modeling the spike. We estimated the gradient along the diffraction spike assuming that it can be modeled by one polynomial function that applies to each row of pixels. To obtain this polynomial function, we calculated the -clipped median of each column across the width of the spike: 150 pixels for the eastern spike of Merope, 100 pixels for the other spikes. This gave us the typical longitudinal gradient profile of the spike, which a fit with a seventh-order polynomial function (see Fig. 13).
@@
Finally, we normalized the first approximation of the spike we obtained earlier by dividing it, for each column across the spike, by the polynomial values of the longitudinal gradient. For each column, we then selected the nearest 100 columns to compute the -clipped median of this selection for each row. This gave us the profiles across the spike, assuming it does not vary on scales much shorter than 100 pixels or 6.3″(see Fig. 14).
@@
We note that the pattern varies smoothly along the spike. In particular, we see the peak of the profile shifts by about 2 pixels over 2000 pixels. A larger shift seems to occur at the faintest end of the spike, far from Merope, where the relative intensity of the spike is the lowest. However, even there, the shift is not greater than 7 pixels compared to the other end of the spike. Next, we rescaled the normalized profiles with the polynomial gradient to obtain the spike model shown in the fourth panel of Fig. 12. Finally, we subtracted the model from the initial map (see the fifth panel in Fig. 12). To assess the accuracy with which the spike is removed, we compared the residual background in the spike image, away from the central area, relative to the initial map. We find that, in the case of Merope’s straight spike, the residual is typically lower than 1% at more than 75 pixels (4.7 ″) from the center of the spike (see Fig. 15).
@@
A limitation in our method is visible in Fig. 16. Closer to Merope, the pattern across the spike varies on scales smaller than 100 pixels (or 6.1″), leaving a residual trace near the edge of our observations. However, the alternative (i.e., using a smaller window than 100 pixels) would not properly filter out the sources “hidden” by the spikes. We compared these residuals with local dust emission by normalizing the profiles at the site of the spike to their median values (see Fig. 17). In most cases, the residual of the spike is barely noticeable above the intrinsic variations of the diffuse emission, at a level of 5% of the median value of the local background. However, near Merope, the residual reaches more than 10% of that level but remains mostly random.
@@
At this point, we thus have four mosaics, two for each field, with all large diffraction spikes removed. For each pair of mosaics, we removed the glitches by comparing the two mosaics. Each pixel whose value varies by more than 0.5 MJy/sr was replaced by the lower value in the pair.
A.2 Multi-scale background
@@
Once the large spikes have been removed from the data, we modeled the diffuse emission. We used mosaics of Fields 1 and 2 from which the spikes have been removed. We computed their global median by clipping pixels above 2 and subtracted the result from the images.
@@
Next, we subtracted the different scales of the diffuse emission by repeating three times the following steps: (1) we computed the standard deviation of the 2-clipped background subtracted image, (2) we masked the bright pixels whose values are more than 15 above the median of the entire image, (3) we performed a 2D median-filtering with a 4 clipping without the masked pixels, and (4) we subtracted the median-filtered image. We used the Background2D class from photutils to perform these steps (Bradley et al. 2023). We applied the 2D median filter using rectangular boxes elongated along the direction of the detector arrays, which is close to the orientation of the striations. We used boxes with mean sizes of 100, 40, and 25 pixels. The ratio between the long and short sizes is 2 at all three scales. Figure 18 illustrates the composition of the background map as smaller scales are included, as well as the residual map obtained after background subtraction for a sub-image of Field 1. There is structure in the diffuse emission at scales smaller than 25 pixels (i.e., 1.5 arcsec). This contributes noticeably, albeit minimally, to the residuals, even in Field 1, where the diffuse emission is strongest.
@@ A.3 Removing point sources within the field
After subtracting the multi-scale background, we used the residual images to identify point sources and model the PSF. We identified point sources by recursively selecting the brightest pixel in the four combined mosaics. For each such pixel, we looked for all other neighbor bright pixels and labeled the cluster as a source before looking for the next brightest pixel in the observations. We extracted in this way the brightest 400 sources from our observations. We then visually inspected them all to separate point sources from extended sources, and ended up with 175 point sources. We did not use publicly available source extraction algorithms as they tend to identify multiple sources within each JWST PSF, as those are highly structured for the brightest sources. Our method is more robust and still functional given the small number of sources we wanted to extract.
To build an empirical PSF (ePSF), we followed the method developed by Anderson and King (2000) and Anderson (2016), which is implemented as the EPSFBuilder class in photutils, adapting it to our specific needs. In particular, we wanted to clean the spikes of the point sources more thoroughly than their cores because the spikes contaminate the diffuse emission more than the cores.
To combine the images of the selected point sources, we first centered each image to sub-pixel accuracy using the theoretical PSF of WebbPSF (Perrin et al. 2014) on a 301 by 301 pixel ( by ). Each PSF is fitted in a squared annulus region, instead of its core, and we find that using a region of 45 by 45 pixels, with the 9 by 9 innermost pixels masked, gives the best results. The difference between the position of the brightest pixel and the computed centers of the PSFs is typically less than 0.4 pixel.
The centered PSFs are then stacked in an iterative way by our modified EPSFBuilder to generate the ePSF for our observations. Here, we used the same square annulus as before to align the individual PSFs with each other. The method also returns the individual best fit for each source in our collection, in their 301 by 301 original pixel grid, which we subtracted from the observations, except for the six brightest point sources that have diffraction spikes extending beyond 20″ (see Fig. 19). There are four point sources in Field 1 and two in Field 2, for which the core is highly saturated. If we had removed their ePSF best fit, this would have led to significant residuals across the entire 301 by 301 pixel region. For these sources, we instead waited until we added back the multi-scale background, after which we removed their diffraction spikes in the same way as for the stars off field (see Appendix A.1).
@@ A.4 Masking the remaining sources
At this stage, we had removed all diffraction spikes from sources within and without our field of observations. We were left with masking the saturated core of bright point sources and compact sources.
To build the mask of the remaining sources, we first identified pixels that correspond to significant residuals after subtracting the ePSF fit. We computed the 2-clipped standard deviation for each 301 by 301 pixel image of a point source, keeping only values above a floor of 0.1 MJy/sr. Any pixel in the image whose value in the residual of the ePSF fit is greater than 10 times this standard deviation is flagged. Any pixel whose value in the ePSF fit is greater than 15 times this standard deviation is also flagged. These pixels represent fewer than 0.5% of all pixels in the four sub-mosaics and are the first included in a mask.
Next, we masked additional bright pixels. We started by creating a multi-scale background in an identical manner to what was done in the previous subsection, but now starting from a cleaner map. We then selected all pixels where the residual emission after subtraction of the multi-scale background is higher than a brightness threshold (0.035 MJy/sr for the fiducial mask). We then identified small and large clusters of masked pixels: those with fewer than or more than nine connected masked pixels. We verified that the small clusters typically correspond to faint stars or isolated glitches. Large clusters correspond either to the core of galaxies, whose shapes and sizes are very diverse (see Fig. 11), or to the few bright point sources for which the ePSF method was not applied. Since the brightness threshold leaves the edges of bright sources unmasked, we extended the large clusters by a buffer zone of 15 additional pixels. To do this, we used the skimage.morphology package (van der Walt et al. 2014). Finally, we added back the small clusters of pixels to the mask.
Starting from mosaics containing large diffraction spikes, point sources, and extended sources, we removed the spikes step by step from both the off-field and in-field sources and masked both the core of the point sources and the extended sources. This left us with mosaics containing only diffuse emission, which were used to perform the Fourier analysis presented in Sect. 3.
@@ Appendix B Magnetized interstellar matter near the Pleiades stars
We used the dust temperature and opacity maps derived from fitting the dust spectral energy distribution at far-infrared wavelengths to a modified black body (Planck Collaboration 2016b), and the dust polarization maps at 353 GHz (Planck Collaboration 2020b). We also used the full sky image produced by Meisner and Finkbeiner (2014) from the WISE data at m151515Image downloaded from https://faun.rc.fas.harvard.edu/ameisner/wssa/healpix.html. The angular resolution of these maps is 5’. However, Stokes and are smoothed to 7’ in order to improve the signal-to-noise ratio of the dust polarization.
The Planck and WISE maps offer complementary insights into the ISM. The dust temperature traces the localized heating of dust grains by the Pleiades stars, while the dust opacity indicates the gas column density. The conversion factor from to for translucent molecular gas is (Planck Collaboration 2014). The dust polarization angles are perpendicular to the magnetic field component in the plane of the sky. The WISE image traces the emission of PAHs, the brightness of which scales linearly with the product of the intensity of the radiation field and the gas column density (Draine and Li 2007; Compiègne et al. 2011). Measurements of these observational tracers computed at the center positions of the four JWST images and for the surrounding background are listed in Table 7. The background measurements correspond to the median value calculated for pixels within of Merope where is smaller than to exclude the Pleiades area.
Figure 20 shows the Planck maps of dust temperature and submillimeter opacity map at GHz overlaid on the m WISE image at 12″resolution161616The image was downloaded from https://irsa.ipac.caltech.edu/applications/wise. At the distance of the Pleiades stars, the angular size of the image corresponds to 3 pc. The dust temperature exceeds the background value across the full extent of the mid-infrared emission from the nebula. The opacity contours delineate an interstellar cloud to the south of Merope. It corresponds to a small cloud identified by Ungerechts and Thaddeus (1987) in their survey with a peak CO(1-0) brightness of . The Merope star and the two JWST fields are located at the edge of this translucent cloud. The overlap between the contours of and indicates that the cloud is locally heated by Merope.
@@ Appendix C Association with the cold neutral medium
In the paper introduction, we claim that the JWST Pleiades observations probe the structure of the CNM within a few parsecs of Merope. To support and discuss this statement, we used the Planck and WISE data presented in Appendix B. These data provide a global perspective on the JWST observations. The angular resolution of the Planck data corresponds to a scale of 0.2 pc at the distance of Merope.
@@
Merope lies to the front of the large (150 pc in diameter) Per-Tau bubble that was identified by Bialy et al. (2021) using 3D stellar reddening data. The star is close to the outer surface of the shell that surrounds the bubble. On smaller scales, the dust emission data presented in Appendix B show that Merope and the two JWST fields are located at the edge of a translucent cloud that has a diameter of a few parsecs. Table 7 lists the values of the dust temperature and the dust opacity in the center positions of the four JWST images, as well as for the surrounding background. The values of the Planck Stokes parameters at 353 GHz, and the PAH emission at m, measured on the WISE map of the entire sky produced by Meisner and Finkbeiner (2014), are also listed in the table. We used these data to estimate the radiation field intensity and gas column density at the location of the JWST fields.
@@
The radiation field intensity can be estimated from the dust temperature using the formula
@@
with a spectral index (Planck Collaboration 2014). The factor measures the radiation field intensity normalized to its value within the background ISM, away from the Pleiades stars. This formula yields values ranging from 6 to 9 at the position of the JWST fields. The opacity values listed in Table 7 include the Galactic background. To estimate the column density of the cloud associated with Merope, we needed to subtract this contribution. This subtraction introduces a large uncertainty equal to the dispersion of opacity values measured over the background area, which is listed as the error bar in the background line of the table. Taking into account this uncertainty, we assumed that the column density of the cloud at the positions of the JWST fields cannot be much larger than .
@@
The m brightness allows us to crosscheck the estimates of and . In the solar neighborhood, the mean emission at m is for a column density of (Compiègne et al. 2011). For fixed PAH abundance and emission properties, this emission scales linearly with the product (Draine and Li 2007; Compiègne et al. 2011). Based on the m brightness values in Table 7, we estimated the product to be 16 and 7 for Fields 1 and 2, where is expressed in units of . These values suggest that the estimates of G and derived from dust temperature and opacity are somewhat underestimated. This could be due to beam smoothing and the superposition of cloud and background emissions, which were not considered when fitting the dust temperature and opacity. In any case, the estimates of and support our assertion that the JWST observations characterize emission from CNM within a few parsecs of Merope. The value of is 10 at a distance of 0.5 pc from Merope. The contribution to of warm gas with densities of a few tenths of a across this distance is negligible.
@@
We also estimated the contribution of the Galactic background to the JWST power spectra. To achieve this, we used the scaling law established by Miville-Deschênes et al. (2007) when analyzing Infrared Astronomical satellite (IRAS) data. The fraction of power contributed by the Galactic background is estimated using the following formula:
@@
where , the brightness in the center of a JWST image in the F335M filter, is obtained by multiplying the values and in Table 1. To obtain the background brightness in the F335M filter, we scaled the background brightness at m listed in Table 7 by the intensity ratio , as determined in the study of Flagey et al. (2006). We obtain values of % for the Field 1 images and between one and a few percent for the Field 2 images. These low values support the link that the paper makes between the JWST power spectra and the CNM near the Pleiades stars.